### v1

@@ Abstract
NL2VIS (natural language to visualization) is a promising and recent research area that involves interpreting natural language queries and translating them into visualizations that accurately represent the underlying data. As we navigate the era of big data, NL2VIS holds considerable application potential since it greatly facilitates data exploration by non-expert users. Following the increasingly widespread usage of generative AI in NL2VIS applications, in this paper we present V-RECS, the first LLM-based Visual Recommender augmented with explanations (E), captioning (C), and suggestions (S) for further data exploration. V-RECS’ visualization narratives facilitate both response verification and data exploration by non-expert users. Furthermore, our proposed solution mitigates computational, controllability, and cost issues associated with using powerful LLMs by leveraging a methodology to effectively fine-tune small models, like LLama-2-7B. To generate insightful visualization narratives, we use Chain-of-Thoughts (CoT), a prompt engineering technique to help LLM identify and generate the logical steps to produce a correct answer. Since CoT is reported to perform poorly with small LLMs, we adopted a strategy in which a large LLM (GPT-4), acting as a Teacher, generates CoT-based instructions to fine-tune a small model, Llama-2-7B, which plays the role of a Student. Extensive experiments - based on a framework for the quantitative evaluation of AI-based visualizations and on manual assessment by a group of participants - show that V-RECS achieves performance scores comparable to GPT-4, at a much lower cost. The efficacy of the V-RECS teacher-student paradigm is also demonstrated by the fact that the un-tuned Llama-2-7B fails to perform the task in the vast majority of test cases. We release V-RECS for the visualization community to assist visualization designers throughout the entire visualization generation process.
keywords:
Visualization recommendation, LLM, Machine Learning for Visualization, Chain of Thoughts, Low-cost Generative AI
[…]
During the training phase, a large LLM, GPT-4, acting as a Teacher (B), receives as input triples (D,Q,V) extracted from an available dataset, where D is tabular data, Q is a natural language query, and V is the VegaZero specification for a chart matching the user’s query Q. The model is asked to perform three Chain of Thought (CoT) tasks (A): T1, to explain the reasoning steps that justify the answer V given the query Q; T2, to generate a caption that describes the chart, and T3, to suggest additional useful queries. The responses of the model are combined to create a visualization narrative, VN (C,D). Next, a small model, Llama-2-7B (the Student), is fine-tuned (E) with the enriched dataset of quadruples (D,Q,C,VN). At inference time, the resulting specialized model, named V-RECS, receives as input a pair (D,Q) and generates, along with V, a narrative made of an explanation E, a caption C, and a suggestion S of additional queries. Adopting a teacher-student metaphor, the esteemed professor GPT-4, leveraging its trillion-parameters knowledge of data visualization and natural language understanding, generates examples of visualization narratives for the young student Llama-2-7B. Llama uses this supplementary material to diligently learn the task of producing similar narratives alongside the recommended visualization, and finally gets its doctorate in Visual Recommendation with Narratives, becoming Dr. V-RECS. During the exercise of its profession, will Dr. V-RECS be able to match Professor GPT-4’s abilities? Certainly, its fees are much cheaper.
Introduction
1 Introduction
@@
AI and Big Data development has reshaped how companies make decisions in recent years. A more data-driven approach has become increasingly crucial for companies to make more informed decisions, from big corporations to small and medium enterprises [49]. From marketing [40] to business intelligence, the data-centric strategy represents a terrific opportunity to keep up with a specific field’s evolution and discover patterns that otherwise would be hidden. Notwithstanding, this introduces new challenges and complexity. The data analysis and transformation into actionable insights demand specific domain experts and economic resources. Often, especially for SMEs, these factors represent challenging aspects due to the need for more financial resources to recruit these specialized profiles. Furthermore, the supply of these specialized data scientists often falls short of meeting the burgeoning needs of the industry [^1] https://shorturl.at/ePUXY.
@@
Visualization Recommendation Systems (VRSs) try to fill the gap between non-expert users and data analysis [36], by providing automatic tools to support visual discovery from data across different sectors and applications. Many contributions have been developed in this field (among others, [15, 26, 38, 29]), some of which gained widespread adoption in well-known tools such as Tableau222www.tableau.com, which integrates a functionality named Suggest Me [30]. To date, state-of-the-art VRSs do not achieve full automation, relying on built-in knowledge (such as perceptual rules) or user input to guide the system to generate visualizations, for example by entering natural language queries. This latter category of VRS is referred to in the literature as NL2VIS (Natural Language to Visualization). Many NL2VIS solutions rely on classic NLP (Natural Language Processing) or rule-based techniques [42, 13, 32] to translate the user’s query, called Natural Language Query (NLQ), into insightful visualizations. However, reliance on these techniques poses a significant limitation, as it constrains users to adhere to specific formats (such as the need to clearly specify in the NLQ the type of chart and the variables to be plotted on the axes), thereby restricting the flexibility of expression in their queries. Furthermore, these approaches struggle to manage ambiguous texts, failing to produce a consistent and insightful visualization in many cases.
@@
Following the disruptive use of generative artificial intelligence across application domains, LLM4VIS (Large Language Model for Visualizations) [53] is becoming one of the most promising research areas to advance existing NL2VIS approaches. LLM4VIS aims to use cutting-edge LLM technology to query data sources using only natural language queries and translate them into visualizations that accurately represent the underlying data. Although LLM like Llama [51] or GPT [34], have shown a more advanced semantic understanding of free text [7], they suffer from the lack of interpretability of the model’s responses. Furthermore ”Large" LLM, those involving computationally intensive and expensive architectures containing billions, and recently trillions333as postulated by several sources, among which https://the-decoder.com/gpt-4-has-a-trillion-parameters/ of parameters, also have other contraindications, such as hallucinations [16], the lack of controllability [47] and interpretability [54], and the high computational cost444in https://shorturl.at/puz24 -among other published experiments - it is shown that Llama 2 obtains similar performance to GPT-4 for a fraction of the cost on a number of tasks, hindering their large-scale adoption in applications. In particular, controllability has a notable, and often neglected, impact on applications: if a certain model is based on the adaptation of a proprietary LLM, such as GPT, the evolution of the original LLM [^5] For example, there are notable differences between GPT-3 and GTP-4. could significantly impact the adapted model. This means that developers of a model based on a proprietary LLM do not have full control of their system.
@@
Finally, although LLMs have proven great capabilities on tasks like text, images, or code generation, it has been observed that they are much less proficient on other tasks, such as data visualization and analysis. In this regard, Podo et al. [35] observed how, even though GPT-3.5-turbo can already perform data visualization tasks, it still produces a high rate of errors and poses a problem of response verification, especially when the user is not sufficiently expert to evaluate the result. Unlike text and images, which can be easily verified by any user, or code, which can be verified by compiling and then validating the expected output, evaluating the quality of a recommended visualization is more challenging. If users do not know what to expect, or they are not domain experts, they may perceive the model’s output as accurate, even when incorrect.
@@
To mitigate some of the outlined problems, in this paper we propose V-RECS, a LLM-based Visualization Recommender enhanced with Explanations (E), Captioning (C), and Suggestions for further data exploration (S). The data visualization narratives E and C mitigate the problem of response verification, while S support non-expert users to further explore the available data. The generated model, based on a fine-tuned ”small" LLM, Llama-2-7B, also mitigates cost and controllability issues [^6] Llama-2-7B has seven billion parameters, while, e.g., GPT-4 has one trillion. Furthermore, Llama-2-7B is available free of charge for research and commercial use; in contrast, GPT-n models are proprietary, and their inner workings are hidden from the public https://shorturl.at/bdoOZ.
@@
V-RECS’ training methodology is based on Chain of Thoughts (CoT), a strategy to instruct a large LLM to break down its responses into reasoning steps[55]. Since CoT has been shown not to work well with small LLM [53], we adopt a teacher-student paradigm in which GPT-4, the teacher, is asked to enrich with visualization narratives an existing dataset of users’ queries and corresponding charts, leveraging the CoT strategy. Next, the enriched dataset is used to fine-tune a small open-source LLM, Llama-2-7B, to perform visualization recommendations along with textual narratives to explain and further explore the input data. V-RECS exploits the synergies between LLM and visualization to improve the generation of insightful visual representations from data, while ensuring that the user has full control over the reliability of what is suggested by the Visual Recommender. The main contributions of this work are:
@@
  • V-RECS, the first LLM fine-tuned for the visualization field that performs visualization recommendation along with the production of insightful visualization narratives;
@@
  • Low-cost and controllable solution: V-RECS is based on fine-tuning the open source Llama-2-7B, which was estimated to be 100 times less expensive than GPT-4, along with the other previously summarized advantages of low-parameter Open-AI models 777amongst the many studies, see e.g.,https://tinyurl.com/56kpjd7r. Despite leveraging a small LLM, V-RECS achieves performances comparable to GPT-4, as shown in Section 6.
@@
  • LLM4VIS methodology: we present an end-to-end solution for LLM4VIS, contributing a novel approach that bridges the gap between generative artificial intelligence, explainable machine learning, and the visualization field.
@@
  • A comprehensive comparative quantitative and qualitative evaluation experiment, to study how the proposed V-RECS model and GPT-4 perform the visualization task.
We release V-RECS for the visualization community on a GitHub repository [^8] https://github.com/lucapodo/V-RECS.git.
@@
The paper is organized as follows: in Section 2 we summarize the main concepts of LLM. Section 3 is dedicated to the state-of-the-art in NL2VIS and LLM4VIS. Section 4 presents a formal description of the LLM4VIS problem. Section 5 describes the proposed V-RECS model and the LLM-based methodology to develop it. Section 6 reports the results of quantitative and qualitative evaluation activities. Finally, we present a discussion and our concluding remarks in Section 7.
2 Background
@@
LLMs are powerful machine learning models based on the concept of pre-training followed by fine-tuning. During pre-training, the model learns representations of language through unsupervised learning on extremely vast corpora of textual data. This phase is crucial for the acquisition of general linguistic knowledge. Pre-training is based on Transformer-based architectures [23], capable of capturing long-range dependencies and contexts, essential for understanding the complexities of human language. Following pre-training, fine-tuning tailors the LLM to specific downstream tasks by exposing it to labeled data and adjusting its parameters through supervised learning. This process adapts the model’s learned representations to the specific task at hand, enhancing its performance. Recently, LLMs have been adapted to process and generate different types of unstructured data beyond text, such as images and videos, music, climate data, code, and visualizations, which is the main focus of this paper.
@@
3 Related work
@@
A Visualization Recommendation System (VRS) is a system designed to suggest insightful visualizations given a dataset. As outlined by Podo et al. [36], there are two classes of VRS: task-agnostic and task-aware. The task-agnostics [15, 11] require solely the dataset as input without additional guidance to generate insightful visualizations. Most of the approaches rely on classical machine or deep learning methods. Recent advancements have explored leveraging LLMs [10, 53] to perform this task. In this case, the visualization recommendation process is split into multiple steps (e.g., dataset filtering, dataset description, column selection), with the LLM involved in all or some of these steps.
While attaining full automation in visualization recommendation pipelines would signify a remarkable milestone, it is still in the early stages, facing several challenges that have not yet been explored [36]. Conversely, task-aware VRS requires user input or guidance, typically as user queries, to generate visualizations. Task-aware VRSs are extensively studied in the literature and fall within the domain of Natural Language to Visualization (NL2VIS). In NL2VIS, the system is guided by a natural language query that tells the model what to explore in a dataset (e.g.: What are the payment date of the payment with the amount paid higher than 300 or with payment type is check, and count them by a bar chart, order by the y axis from low to high?). Recently, the field has been significantly influenced by LLMs, which demonstrate considerable ability in understanding human language, allowing them to address the challenge of ambiguity inherent in user expressions that currently limit existing approaches.
3.1 NL2VIS approaches
The development of NL2VIS has witnessed significant advancements, primarily through classical NL techniques.
As discussed by Setlur et al. [42], the authors introduce an interactive visualization tool that employs a hybrid approach integrating NLP and decision rules. This system combines a probabilistic grammar approach with predefined rules that can be dynamically updated based on the underlying data. Another example is proposed by Srinivasan et al. [45], where the authors present BOLT, a system based on traditional NLP techniques to map user queries to prevalent dashboard objectives and generate appropriate visualizations. In contrast, the solution proposed by Narechania et al. [33] focuses on generating a single visualization rather than suggesting a complete dashboard.
While these approaches provide reliable solutions, they impose input constraints by asking users to adhere to specific input templates rather than freely using natural language (for example, explicitly naming the variables to be considered in the chart, avoiding semantically ambiguous sentences, etc.). A departure from these syntactic approaches is NcNet, proposed by  [28]. The authors propose to use a seq2seq neural network model to translate user queries into visualizations, leveraging a novel visualization grammar named VegaZero. VegaZero streamlines the complex Vega-Lite specification into a more intuitive sequence-based keyword format. Generating a plain text sequence proves to be more manageable for a seq2seq model, in contrast to the challenges associated with hierarchical outputs, such as for Vega-Lite.
Despite the advancements introduced, classic NLP systems encounter various challenges. Primarily, they depend on the well-structured syntax of user queries for effective processing. Furthermore, failure to understand the semantics of the query fails to understand what the user wants to explore.
3.2 Large Language Models for visualization
@@
Within the field of NL2VIS, Large Language Models for Visualization (LLM4Vis) is emerging as the new frontier for translating user queries into consistent visualizations. LLM4VIS seeks to integrate LLMs into data visualization pipelines to achieve full support for visualization automation. They have shown an increasing proficiency in comprehending the underlying semantics of human instructions [14], representing a compelling opportunity to overcome the limitations inherited by the ”classic" approaches illustrated in the previous section. While this technology is still in its early stages and requires further exploration and study, it nonetheless presents an emerging opportunity for the visualization field, as discussed by Schetinger et al. [41]. We can categorize the LLM4VIS approaches into two main classes. The first class encompasses all the methods that involve the model in its standard configuration (the so-called ”plain foundation models") and use prompt engineering techniques to improve the model’s performance. Conversely, the second class includes all the approaches that involve a "fine-tuning" phase to specialize the foundation model toward a specific set of tasks.

### v2

@@ Abstract
NL2VIS (natural language to visualization) is a promising and recent research area that involves interpreting natural language queries and translating them into visualizations that accurately represent the underlying data. As we navigate the era of big data, NL2VIS holds considerable application potential since it greatly facilitates data exploration by non-expert users. Following the increasingly widespread usage of generative AI in NL2VIS applications, in this paper we present V-RECS, the first LLM-based Visual Recommender augmented with explanations (E), captioning (C), and suggestions (S) for further data exploration. V-RECS’ visualization narratives facilitate both response verification and data exploration by non-expert users. Furthermore, our proposed solution mitigates computational, controllability, and cost issues associated with using powerful LLMs by leveraging a methodology to effectively fine-tune small models, like LLama-2-7B. To generate insightful visualization narratives, we use Chain-of-Thoughts (CoT), a prompt engineering technique to help LLM identify and generate the logical steps to produce a correct answer. Since CoT is reported to perform poorly with small LLMs, we adopted a strategy in which a large LLM (GPT-4), acting as a Teacher, generates CoT-based instructions to fine-tune a small model, Llama-2-7B, which plays the role of a Student. Extensive experiments - based on a framework for the quantitative evaluation of AI-based visualizations and on a manual assessment by a group of participants - show that V-RECS achieves performance scores comparable to GPT-4 at a much lower cost. We release V-RECS for the visualization community to assist visualization designers throughout the entire visualization generation process.
Index Terms:
Visualization recommendation, LLM, Machine Learning for Visualization, Chain of Thoughts, Generative AI
I Introduction
@@
AI and Big Data development has reshaped how companies make decisions in recent years. A more data-driven approach has become increasingly crucial for companies to make informed decisions [1, 2]. This introduces new challenges and complexity. The data analysis and transformation into actionable insights demand specific domain experts and economic resources. Furthermore, the supply of these specialized data scientists often falls short of meeting the needs of the industry [^1] https://shorturl.at/ePUXY.
@@
Visualization Recommendation Systems (VRSs) try to fill the gap between non-expert users and data analysis [3] by providing automatic tools to support visual discovery from data across different sectors and applications. Many contributions have been developed in this field (e.g., [4, 5, 6, 7]), some of which gained widespread adoption in well-known tools such as Tableau222www.tableau.com integrates a functionality named Suggest Me [8]. To date, state-of-the-art VRSs do not achieve full automation, relying on built-in knowledge (such as perceptual rules) or user input to guide the system in generating visualizations. Although fully automated solutions may lack the flexibility for detailed customization that expert users might require, they can also help overcome several limitations of existing approaches:
(i) Human visual perception, including color perception, pattern recognition, gestalt principles (e.g., proximity, similarity), cognitive load, and more. Capturing these processes’ complexity and their interactions in a comprehensive set of rules is extremely challenging. Furthermore, rule-based approaches require manual updates to accommodate new scenarios. Recent advancements in deep learning have enabled visual recommender systems to automatically learn perceptual rules from data/visualization pairs [3].
(ii) Required expertise for integrating user specifications in the VRSs, making it more responsive and aligned with individual user preferences. On the other hand, this interaction implies knowledgeable users. Users without a data analysis or visualization background may find it challenging to provide the correct query, indicate the right chart type, or accurately interpret the data. NL2Vis (Natural Language to Visualization) systems [9, 10, 11] are being developed to simplify the interaction of non-expert users with complex data, by enabling them to generate visualizations through natural language queries (NLQ). However, reliance on these techniques poses a significant limitation, as it constrains users to adhere to specific formats (such as the need to clearly specify in the NLQ the type of chart and the variables to be plotted on the axes), thereby restricting the flexibility of expression in their queries. Furthermore, these approaches struggle to manage ambiguous texts, failing to produce a consistent and insightful visualization in many cases, as reported, for example, by Luo et al. in their work [12].
A further significant step forward in the automation of visual recommender systems is represented by LLM4VIS (Large Language Models for Visualization) [13]. By leveraging generative AI in conjunction with visualization techniques, LLM4VIS allows it to query data sources and translate them into visualizations that accurately represent the underlying data, showing good capabilities in understanding and processing complex natural language instructions. It represents a promising direction in data visualization, offering advantages in terms of simplicity of user interaction, scalability, and ability to handle complex queries. However, even for these systems, many challenges remain open:
  • Interpretability and Trust: Although LLM like Llama [14] or GPT [15] have shown a more advanced semantic understanding of free text [16], they suffer from the lack of interpretability of the model’s responses [17] and hallucinations [18].
@@
  • Controllability and cost: “Large” LLMs, those involving computationally intensive and expensive architectures containing billions, and recently trillions of parameters, also have other contraindications, such as the lack of controllability [19] and the high computational cost333in https://shorturl.at/puz24 it is shown that Llama 2 obtains similar performance to GPT-4 for a fraction of the cost on a number of tasks, hindering their large-scale adoption in applications. In particular, controllability has a notable impact on applications: if a certain model is based on the adaptation of a proprietary LLM, such as GPT, the evolution of the original LLM could significantly impact the adapted model. This means that developers of a model based on a proprietary LLM do not have full control of their system.
@@
  • Adaptability and Verifiability: Finally, although LLMs have proven great capabilities on tasks like text, images, or code generation, it has been observed that they are much less proficient on other tasks, such as data analysis and visualization. In this regard, Podo et al. [20] observed how, even though GPT-3.5-turbo can already perform data visualization tasks, it still produces a high rate of errors and poses a problem of response verification, especially when the user is not sufficiently expert to evaluate the result. Evaluating the quality of a recommended visualization is a challenging task, and if users do not know what to expect, or they are not domain experts, they may perceive the model’s output as accurate, even when incorrect.
@@
The combination of these issues can lead to situations where the user may undertrust or overtrust the model’s generated output. To better support AI-based visualization generation while addressing its reported issues, in this paper we propose V-RECS, a LLM-based Visualization Recommender enhanced with three types of narratives: Explanations (E), Captioning (C), and Suggestions for further data exploration (S). While E and C mitigate the problem of response verification, better supporting the trustability of the generation process for a user, S supports non-expert users to further explore the data and fulfill data analysis tasks. The generated model, based on a fine-tuned “small” LLM, Llama-2-7B, also mitigates cost and controllability issues [^4] Llama-2-7B has seven billion parameters. Furthermore, Llama-2-7B is available free of charge for research and commercial use.
@@
To achieve this, we also contribute V-RECS’ training methodology. It is based on Chain of Thoughts (CoT), a strategy to instruct a large LLM to break down its responses into reasoning steps [21]. Since CoT has been shown not to work well with small LLM [13], we adopt a teacher-student paradigm in which GPT-4, the teacher, is asked to enrich with visualization narratives an existing dataset of users’ queries and corresponding charts, leveraging the CoT strategy. Next, the enriched dataset is used to fine-tune in a supervised manner a small open-source LLM, Llama-2-7B, to perform visualization recommendations along with producing textual narratives to explain and further explore the input data. V-RECS improves the generation of insightful visual representations from data while ensuring that the user has full control over the reliability of what is suggested by the Visual Recommender.
Summarizing, the main contributions of this work are:
@@
  • V-RECS, the first LLM fine-tuned for the visualization field that performs automated visualization recommendation along with the production of insightful visualization narratives;
@@
  • Low-cost and controllable solution: V-RECS is based on fine-tuning the open-source Llama-2-7B, presenting previously summarized advantages of low-parameter Open-AI models 555amongst the many studies, see e.g.,https://tinyurl.com/56kpjd7r. Despite leveraging a small LLM, V-RECS achieves performances comparable to GPT-4, as shown in Section VI.
@@
  • LLM4VIS methodology: we present an end-to-end solution for LLM4VIS, contributing a novel approach leveraging contributions from generative AI, explainable machine learning, and the visualization field.
@@
  • A comprehensive comparative quantitative and qualitative evaluation experiment to study how V-RECS and GPT-4 perform the visualization recommendation task.
We release V-RECS and its documentation on a GitHub repository [^6] https://github.com/lucapodo/V-RECS.git.
@@
The paper is organized as follows: in Section II we summarize the main concepts of LLM. Section III is dedicated to the state-of-the-art in NL2VIS and LLM4VIS. Section IV presents a formal description of the LLM4VIS problem. Section V describes the proposed V-RECS model and the LLM-based methodology to develop it. Section VI reports the results of quantitative and qualitative evaluation activities. Finally, we present a discussion and our concluding remarks in Section VII.
II Background: LLM training techniques
@@
LLMs are powerful machine learning models based on the concept of pre-training followed by fine-tuning. During pre-training, the model learns representations of language through supervised learning on extremely vast corpora of textual data. This phase is crucial for the acquisition of general linguistic knowledge. Pre-training is based on Transformer-based architectures [22], capable of capturing long-range dependencies and contexts, essential for understanding the complexities of human language. Following pre-training, fine-tuning tailors the LLM to specific downstream tasks by exposing it to labeled data and adjusting its parameters through supervised learning. This process adapts the model’s learned representations to the specific task at hand, enhancing its performance. Recently, LLMs have been adapted to process and generate different types of unstructured data beyond text, such as images and videos, music, climate data, code, and visualizations, which is the main focus of this paper.
@@
Chain of thoughts (CoT) [21] is a prompting technique to generate explanations or to perform complex reasoning tasks. It pushes the model to emulate human-like cognitive processes by telling it to break down complex problems into more manageable sub-problems, enhancing response accuracy. CoT can be classified as zero-shot [24] and few-shot [21]. The few-shot configuration necessitates a set of question-answer pair examples in addition to the new one. In this case, the prompt typically comprises multiple question/reasoning-steps/answer triplets, guiding the model through problem-solving. Unlike the few-shot variant, zero-shot CoT does not provide examples of the problem-solving process but only relies on the model’s embedded knowledge. This configuration requires a specific text formulation within the prompt “Let’s think step by step”, telling the model to split the problem into sequential sub-tasks. For instance, a query like “Problem: how much is 2+2? Let’s think step by step” drives the model to break the question and solve each part individually, culminating in a final response, e.g.: Step 1: Identify the operation to be performed: multiplication; Step 2: Multiply the numbers: 2 × 2 = 4; Step 3: The final result is 4.
In our proposed framework we employ both prompting and supervised fine-tuning to tune our model. As explained in Section V, we adopt a CoT zero-shot prompting strategy to instruct GPT to generate rich visualization narratives VN when given a data table D, a NL query Q, and a visualization specification V. These narratives are used to enrich an existing dataset, which is then used to fine-tune in a supervised manner LLama. We call V-RECS the fine-tuned model. At inference time, when D and Q are provided in input, V-RECS produces a recommendation V along with a narrative VN in zero-shot.
III Related work
@@
A Visualization Recommendation System (VRS) is a system designed to suggest insightful visualizations given a dataset. As outlined by Podo et al. [3], there are two classes of VRS: task-agnostic and task-aware. The task-agnostics [4, 26] require solely the dataset as input without additional guidance to generate insightful visualizations. Conversely, task-aware VRS requires user input or guidance, typically as user queries, to generate visualizations.
While attaining full automation in visualization recommendation pipelines would signify a remarkable milestone, it is still in the early stages, facing several challenges that have not yet been explored [3]. Differently, task-aware VRSs are extensively studied in the literature. Natural Language to Visualization (NL2VIS) systems fall within this category. In NL2VIS, the system is guided by a natural language query that tells the model what to explore in a dataset (e.g.: Which product lines generate the most revenue?). The development of NL2VIS has witnessed significant advancements, primarily through classical NL techniques. Various approaches have been proposed [9, 27, 28], incorporating NLP techniques and optional rule-based systems to recommend visualizations through an interactive visualization tool. While these approaches provide reliable solutions, they impose input constraints by asking users to adhere to specific input templates rather than freely using natural language (for example, explicitly naming the variables to be considered in the chart, avoiding semantically ambiguous sentences, etc.). A departure from these syntactic approaches is explored through more advanced NLP methodologies, such as Transformers, marking the initial step toward the LLM4VIS field. NcNet[12] exemplifies these advancements by utilizing a transformer-based architecture and framing the NL2VIS problem as a translation task. Despite being based on an advanced architecture, due to the limited syntactic and semantic variability of the dataset used for training - compared to an LLM training dataset - the authors note that challenges still persist in handling ambiguous and complex natural language queries. This is a limitation, as not understanding the semantics of the query does not allow the system to understand what the user wants to explore. In contrast, LLMs build on a much broader linguistic knowledge, demonstrating a remarkable ability to understand human language and addressing the challenge of ambiguity inherent in user expressions, which limits current approaches. Moreover, while LLMs can be fine-tuned for specific tasks, such as data visualization with narratives (as proposed in this paper), enhancing the capabilities of classic NLP models would require a system redesign. To the best of our knowledge, none of the ”traditional” NL2Vis approaches already incorporates the ability to produce visualization narratives alongside the recommended visualization.
@@
Within the field of NL2VIS, Large Language Models for Visualization (LLM4Vis) are emerging as the new frontier for translating user queries into consistent visualizations. LLM4VIS seeks to integrate LLMs into data visualization pipelines to achieve full support for visualization automation. They have shown an increasing proficiency in comprehending the underlying semantics of human instructions [29], representing a compelling opportunity to overcome the limitations inherited by the “classic” approaches. While this technology is still in its early stages and requires further exploration and study, it nonetheless presents an emerging opportunity for the visualization field, as discussed by Schetinger et al. [30].
We can categorize the LLM4VIS approaches into two main classes. The first class encompasses all the methods that involve the model in its standard configuration (the so-called “plain foundation models”) and use prompt engineering techniques to improve the model’s performance (Section III-A). Conversely, the second class includes all the approaches that involve a “fine-tuning” phase to specialize the foundation model toward a specific set of tasks (Section III-B). In what follows, we summarize the state of the art (SoA) for each category of systems. For completeness, we also dedicate a section to LLM-based systems capable of generating visualization narratives, but not visualization recommendations (Section III-C).
III-A LLM4VIS prompt tuning approaches